bill has $3.75 in quarters,dimes,and nickels. he has twice as many quarters as dimes and three more dimes than nickels. find how much of each coin bill has.
Answer by rossiv53(27)

Let x = the number of nickels, x+3 = the number of dimes, and 2(x+3) = the number of quarters.
So, your equation is
5x+10(x+3)+25(2(x+3))=375
5x+10x+30+50x+150=375
65x+180=375
65x+180-180=375-180
65x=195
x=3
So number of nickels = 3 or .15
Dimes = 6 or .60
Quarters = 12 or 3.00
Value of coins = $3.75
Good luck!